OpenStudy (anonymous):

a power model is shown below. determine the residual for the observed data x=7 and y=70. log y^hat = 1.6 + 0.3 log x a) 71.37 b) 1.37 c)1.85 d) -1.37

OpenStudy (btaylor):

The expected value is what you get when you plug in x. Here, you get log(y-hat)=1.6 + .3 log (7) = 1.85... To get y-hat, we take 10 to this number, so 10^1.85...= 71.37 The residual is the difference between the expected and actual values. So, resid = y - yhat resid = 71.37 - 70 = 1.37 B is your answer.